Calculate the final temperature that will be obtained when a 10g portion of KCL is dissolved in a coffee-cup calorimeter containing 100g of water. Specific heat of the solution is 4.084j/gc

delta H sol (KCL)= 17.0kj/mol and initial temperature is 21.0. Assume that density of the solution is 1.00 g/ml and the calorimeter loses only a neglible amount of heat.
Answer #1

mass of the solution = mass og KCl+ mass of water= 10+100= 110 gm

moles of KCl= mass/molar mass= 10/74.5 =0.134,   1 mole of KCl gives rise to enthallpy change of 17 Kj

0.134 moles of KCl gives rise to enthalpy change= 17*0.134 KJ=2.3 Kj= 2.3*1000 Joules= 2300 joules

Enthalpy change is +ve suggesting that the reaction is endothermic reaction and there is drop in temperature.

2900 = Mass of solution* specific heat of solution* ( 21-T)

21-T= 2900/(110*4.084), T= 14.54 deg.c